OpenStudy (anonymous):

What is the equation of a line which passes through (-2, -3) and has a slope of 1/2?

OpenStudy (anonymous):

Use the point slope form first, and substitute in the values that are given to you. y - y1 = m (x-x1) y + 3 = 1/2 (x +2) NOW, we can just simplify this equation so that it's in slope intercept form. You would distribute out the brackets first, and then just reorder the terms. y + 3 = 1/2x + 1 y + 3 - 3 = 1/2x + 1-3 y = 1/2x -2 Which is now in slope intercept form. x

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Sorry but the answers are either 2x+y=4 2x-y=4 x+2y=4 x-2y=4